# CriSPO: Multi-Aspect Critique-Suggestion-guided Automatic Prompt Optimization for Text Generation

### v1

@@ 1 Introduction
Recent research has made notable progress in automatic prompt engineering for discriminative tasks, such as text classification (Zhou et al. 2022; Yang et al. 2023; Pryzant et al. 2023; Sordoni et al. 2024). These methods focus on optimizing task prompts for a single metric on a single aspect. The process typically involves instructing an LLM optimizer with a meta-prompt to generate new task prompts based on previously sampled task prompts and their corresponding scores. By iteratively exploring candidates and selecting the task prompt with the highest score, performance on the target metric improves over numerous iterations. However, applying these methods directly to text generation tasks, such as summarization, is sub-optimal due to challenges in obtaining effective optimization signals. Unlike classification tasks, where metrics are straightforward (eg. accuracy), automatic metrics for text generation, like ROUGE (Lin 2004), provides limited guidance for prompt refinement. For example, a lower ROUGE score may result from aspects such as mismatched length, differences in word choice due to formality, or varying writing formats, making it difficult to guide LLMs in prompt modification without fine-grained feedback targeting these individual aspects. Furthermore, evaluating text generation involves multiple metrics (Fabbri et al. 2021; Gao and Wan 2022; Elangovan et al. 2024). In addition to reference similarity, other metrics such as factual consistency, which can be assessed using metrics like AlignScore (Zha et al. 2023), is also important. Balancing or utilizing these multiple metrics is not fully addressed by existing prompt engineering methods that focus on optimizing a single metric.
@@
(1) We propose CriSPO, an automatic prompt engineering approach tailored for generative tasks. It discovers aspects to critique generated text and write suggestions for more effective prompt revision.
(2) We conduct comprehensive experiments across multiple LLMs and datasets, demonstrating the effectiveness and robustness of our method. We show an overall 3-4% improvement on ROUGE scores with qualitative verification from human evaluation. CriSPO also obtained consistent improvements on various QA tasks.
@@ 3.1 Multi-Aspect Critiques and Suggestions
$M_{c}$ for different LLMs and tasks are shown in Appendix G.
@@
Table 1: Comparing CriSPO with manual prompts and OPRO on representative summarization benchmarks. Averaged R1/R2/RL (i.e. ROUGE-1/2/L) are reported across 3 runs. 3-shot*: 3-shot ICL with example selection. Claude3: Claude3 Sonnet. Claude In.: Claude Instant. Llama3 results marked with (#) are using Claude3 Sonnet as the optimizer. 3-shot* results are empty for Llama3 due to its limited context window. Standard deviation and SOTA results are in Appendix L. Claude3 Sonnet achieves new SOTA ROUGE-1 performance on D2Note. MBank: MeetingBank; D2Note: Dialogue2Note.
@@ 3.2 Receptive Prompt Optimizer
Specifically, we enhance the OPRO optimizer module with the following three improvements to better utilize critiques and suggestions for achieving stronger guidance and better exploration. See Appendix H for all $M_{o}$ by LLMs and tasks.
@@ 4.1 Experiment Setup
We select a diverse range of 4 summarization tasks including conventional document summarization tasks such as CNN daily mail (Hermann et al. 2015) (news headline summarization), and also conversation summarization tasks such as SAMSum (Gliwa et al. 2019), MeetingBank (Hu et al. 2023). In addition, we test on a medical-domain clinical note summarization task, Dialogue2Note (Yim et al. 2023). Detailed data setup can be found in Appendix C. These tasks cover various lengths, domains and styles as summarized in Table 9. We report ROUGE-1/2/L F-measure (Lin 2004) [^2] We report additional metrics including AlignScore (Zha et al. 2023) and BertScore (Zhang et al. 2019) in Appendix L. to measure output similarity to the references.
@@
Our baseline methods include manual prompts with zero/few-shot ICL. These manual prompts are carefully tuned for each task to incorporate length constraints and task guidelines, and therefore establish a high bar of performance from manual prompt engineering (Appendix I). Given there are no existing automatic prompting results for text generation, we adapted OPRO (Yang et al. 2023), a competitive established approach, on our selected tasks. We use the same hyper-parameter setup in OPRO and CriSPO for fair comparison.
@@ 4.3 Ablating Key Ingredients
The key novelty in our proposed novel critique-suggestion strategy in CriSPO is that it has multi-aspect: i.e. the LLM will generate multi-aspect comparison without enforcing predefined aspects. To understand the effect of the multi-aspect critique-suggestion, we provide two alternative baselines: 1. no multi-aspect: we ask LLM to compare predictions and references in general with no explicit requirement for generating critique and suggestions along multiple dimensions/aspects. This is in line with the approach adopted by Pryzant et al. (2023). 2. predefined aspects: we carefully design dimensions potentially helpful for the summarization task and include verbosity, comprehensiveness, precision and style along with their definitions (Appendix K). The no multi-aspect critique-suggestion baseline performs significantly worse, lacking critical and targeted suggestions due to its tendency to be too general. The predefined multi-aspect approach is as effective as CriSPO but we see no significant improvement from explicit definitions of dimensions. This is because the critique LLM in CriSPO is already able to identify relevant dimensions (such as completeness, verbosity etc. as in Table 8) for each iteration without explicit guidance.
@@ Appendix B Limitations
Dialogue2Note Docter-patient (long) conversation medical note generation 1372 476
@@ Appendix C Dataset Setting
For CNN, SAMSum, MeetingBank, MedMCQA, Narrative QA and SQUAD, we use the HuggingFace datasets repository. For Dialogue2Note, we use the data from Task B at ACL ClinicalNLP MEDIQAChat shared task 2023 in the Acibench dataset [^3] https://github.com/abachaa/MEDIQA-Chat-2023 (Yim et al. 2023). For Natural Questions, we follow the data preparation in FiD [^4] https://github.com/facebookresearch/FiD (Izacard and Grave 2021).
Our experiments are conducted with sampled train and dev set. For Dialogue2Note, we used the full training (67), development (20) and test set (40). For other summarization tasks, we randomly selected 500 samples from the full test set as our test set. To show the efficiency of our approach, we used a small fraction of the train and development set. For CNN, we sampled 100 training samples as our training set, and 100 development samples as our development set. For other tasks, we randomly sampled 50 training samples as our training set, and 50 development samples as our development set.
@@ Appendix F Visualization of Prompt Diversity
Appendix G Multi-Aspect Critique-Suggestion Meta-Prompt
G.1 Claude for Summarization
@@ G.3 Claude for RAG
Appendix H Receptive Optimizer Meta-Prompt
H.1 Claude for Summarization
@@ H.3 Claude for RAG
Appendix I Manual Prompts
@@ I.2 Few-shot CNN
I.3 Zero-shot SAMSum
@@ I.4 Few-shot SAMSum
I.5 Zero-shot MeetingBank
@@ I.6 Few-shot MeetingBank
I.7 Zero-shot Dialogue2Note
@@ I.8 Few-shot Dialogue2Note
I.9 Manual Prompt Tuning
@@
CNN MBank SAMSum D2note
@@
Table 10: Comparing our zero-shot manually tuned prompts (ours) with naive prompts on summarization tasks with Claude Instant. MBank= MeetingBank. D2Note: Dialogue2Note
Appendix J Best QA Prompts Found using CriSPO (Claude Instant)
J.1 Natural Questions
@@
Appendix K Ablation Study Prompts
@@
Appendix L Full Metrics for Summarization
@@
Dataset LLM z-shot 3-shot* OPRO CriSPO CriSPO 3-shot*
@@
D2Note Claude Instant 43.9 51.5 45.2 (±0.2) 53.0 (±0.4) 58.2 (±1.8)
@@
Dataset LLM z-shot 3-shot* OPRO CriSPO CriSPO 3-shot*
@@
D2Note Claude Instant 16.9 23.6 16.3 (±0.4) 19.7 (±0.6) 26.7 (±2.3)
@@
Dataset LLM z-shot 3-shot* OPRO CriSPO CriSPO 3-shot*
@@
D2Note Claude Instant 26.1 33.5 25.5 (±1.0) 26.8 (±1.4) 35.3 (±2.3)
@@
Dataset LLM z-shot 3-shot* OPRO CriSPO CriSPO 3-shot*
@@
D2Note Claude Instant 85.5 88.1 85.1 (±0.3) 85.8 (±0.7) 88.1 (±0.5)
@@
Dataset LLM z-shot 3-shot* OPRO CriSPO CriSPO 3-shot*
@@
D2Note Claude Instant 66.7 66.3 62.3 (±2.0) 63.3 (±3.3) 65.6 (±1.1)
@@
Appendix M Standard Deviation for QA
@@
Dataset Claude z-shot 64-shot OPRO CriSPO CriSPO 64-shot
@@
z-shot 5-shot OPRO CriSPO CriSPO 5-shot

### v3

@@ 1 Introduction
Recent research has made notable progress in automatic prompt engineering for discriminative tasks, such as text classification (Zhou et al. 2022; Yang et al. 2023; Pryzant et al. 2023; Sordoni et al. 2024). These methods focus on optimizing task prompts for a single metric on a single aspect. The process typically involves instructing an LLM optimizer with a meta-prompt to generate new task prompts based on previously sampled task prompts and their corresponding scores. By iteratively exploring candidates and selecting the task prompt with the highest score, performance on the target metric improves over numerous iterations. However, applying these methods directly to text generation tasks, such as summarization, is sub-optimal due to challenges in obtaining effective optimization signals. Unlike classification tasks, where metrics are straightforward (eg. accuracy), automatic metrics for text generation, like ROUGE (Lin 2004), provides limited guidance for prompt refinement. For example, a lower ROUGE score may result from aspects such as mismatched length, differences in word choice due to formality, or varying writing formats, making it difficult to guide LLMs in prompt modification without fine-grained feedback targeting these individual aspects. Furthermore, evaluating text generation involves multiple metrics (Fabbri et al. 2021; Gao and Wan 2022; Elangovan et al. 2024). In addition to reference similarity, other metrics such as factual consistency, which can be assessed using metrics like AlignScore (Zha et al. 2023), are also important. Balancing or utilizing these multiple metrics is not fully addressed by existing prompt engineering methods that focus on optimizing a single metric.
@@
1) We propose CriSPO, an automatic prompt engineering approach tailored for generative tasks. It discovers aspects to critique generated text and write suggestions for more effective prompt revision.
2) We conduct comprehensive experiments across multiple LLMs and datasets, demonstrating the effectiveness and robustness of our method. We show an overall 3-4% improvement on ROUGE scores with qualitative verification from human evaluation. CriSPO also obtained consistent improvements on various QA tasks.
@@ 3.1 Multi-Aspect Critiques and Suggestions
$M_{c}$ for different LLMs and tasks are shown in Appendix H.
@@
Table 1: Comparing CriSPO with manual prompts and OPRO on representative summarization benchmarks. Averaged R1/R2/RL (i.e. ROUGE-1/2/L) are reported across 3 runs. 3-shot*: 3-shot ICL with example selection. Claude3: Claude3 Sonnet. Claude In.: Claude Instant. Llama3 results marked with (#) are using Claude3 Sonnet as the optimizer. 3-shot* results are empty for Llama3 due to its limited context window. Standard deviation and SOTA results are in Appendix M. Claude3 Sonnet achieves new SOTA ROUGE-1 performance on ACI-Bench.
@@ 3.2 Receptive Prompt Optimizer
Specifically, we enhance the OPRO optimizer module with the following three improvements to better utilize critiques and suggestions for achieving stronger guidance and better exploration. See Appendix I for all $M_{o}$ by LLMs and tasks.
@@ 4.1 Experiment Setup
We select a diverse range of 4 summarization tasks including conventional document summarization tasks such as CNN daily mail (Hermann et al. 2015) (news headline summarization), and also conversation summarization tasks such as SAMSum (Gliwa et al. 2019), MeetingBank (Hu et al. 2023). In addition, we test on a medical-domain clinical note summarization task, ACI-Bench (Yim et al. 2023). Detailed data setup can be found in Appendix C. These tasks cover various lengths, domains and styles as summarized in Table 9. We report ROUGE-1/2/L F-measure (Lin 2004) [^2] We report additional metrics including AlignScore (Zha et al. 2023) and BertScore (Zhang et al. 2019) in Appendix M. to measure output similarity to the references.
@@
Our baseline methods include manual prompts with zero/few-shot ICL. These manual prompts are carefully tuned for each task to incorporate length constraints and task guidelines, and therefore establish a high bar of performance from manual prompt engineering (Appendix J). Given there are no existing automatic prompting results for text generation, we adapted OPRO (Yang et al. 2023), a competitive established approach, on our selected tasks. We use the same hyper-parameter setup in OPRO and CriSPO for fair comparison.
@@ 4.3 Ablating Key Ingredients
The key novelty in our proposed novel critique-suggestion strategy in CriSPO is that it has multi-aspect: i.e. the LLM will generate multi-aspect comparison without enforcing predefined aspects. To understand the effect of the multi-aspect critique-suggestion, we provide two alternative baselines: 1. no multi-aspect: we ask LLM to compare predictions and references in general with no explicit requirement for generating critique and suggestions along multiple dimensions/aspects. This is in line with the approach adopted by Pryzant et al. (2023). 2. predefined aspects: we carefully design dimensions potentially helpful for the summarization task and include verbosity, comprehensiveness, precision and style along with their definitions (Appendix L). The no multi-aspect critique-suggestion baseline performs significantly worse, lacking critical and targeted suggestions due to its tendency to be too general. The predefined multi-aspect approach is as effective as CriSPO but we see no significant improvement from explicit definitions of dimensions. This is because the critique LLM in CriSPO is already able to identify relevant dimensions (such as completeness, verbosity etc. as in Table 8) for each iteration without explicit guidance.
@@ Appendix B Limitations
ACI-Bench Docter-patient (long) conversation medical note generation 1372 476
@@ Appendix C Dataset Setting
For CNN, SAMSum, MeetingBank, MedMCQA, Narrative QA and SQUAD, we use the HuggingFace datasets repository. For ACI-Bench, we use the data from Task B at ACL ClinicalNLP MEDIQAChat shared task 2023 in the Acibench dataset [^3] https://github.com/abachaa/MEDIQA-Chat-2023 (Yim et al. 2023). For Natural Questions, we follow the data preparation in FiD [^4] https://github.com/facebookresearch/FiD (Izacard and Grave 2021).
Our experiments are conducted with sampled train and dev set. For ACI-Bench, we used the full training (67), development (20) and test set (40). For other summarization tasks, we randomly selected 500 samples from the full test set as our test set. To show the efficiency of our approach, we used a small fraction of the train and development set. For CNN, we sampled 100 training samples as our training set, and 100 development samples as our development set. For other tasks, we randomly sampled 50 training samples as our training set, and 50 development samples as our development set.
@@ Appendix F Visualization of Prompt Diversity
Appendix G Experiments on Other Tasks
We conduct experiments on other natural language generation (NLG) tasks – DailyDialog (Li et al. 2017) and WebNLG (Gardent et al. 2017). The results are shown on table 10.
Method DailyDialog (R-L) WebNLG (BLEU)
Manual 0-shot 12.6 31.3
Manual 3-shot 17.1 34.8
OPRO 13.3 33.3
CriSPO 17.4 44.3
Table 10: Results on DailyDialog and WebNLG.
Appendix H Multi-Aspect Critique-Suggestion Meta-Prompt
H.1 Claude for Summarization
@@ H.3 Claude for RAG
Appendix I Receptive Optimizer Meta-Prompt
I.1 Claude for Summarization
@@ I.3 Claude for RAG
Appendix J Manual Prompts
@@
J.3 Zero-shot SAMSum
@@
J.5 Zero-shot MeetingBank
@@
J.7 Zero-shot ACI-Bench
@@
J.9 Manual Prompt Tuning
@@
CNN MBank SAMSum ACI-Bench
@@
Table 11: Comparing our zero-shot manually tuned prompts (ours) with naive prompts on summarization tasks with Claude Instant. MBank= MeetingBank.
Appendix K Best QA Prompts Found using CriSPO (Claude Instant)
K.1 Natural Questions
@@
Appendix L Ablation Study Prompts
@@
Appendix M Full Metrics for Summarization
@@
Dataset LLM 0-shot 3-shot* OPRO CriSPO CriSPO 3-shot*
@@
ACI-Bench Claude Instant 43.9 51.5 45.2 (±0.2) 53.0 (±0.4) 58.2 (±1.8)
@@
Dataset LLM 0-shot 3-shot* OPRO CriSPO CriSPO 3-shot*
@@
ACI-Bench Claude Instant 16.9 23.6 16.3 (±0.4) 19.7 (±0.6) 26.7 (±2.3)
@@
Dataset LLM 0-shot 3-shot* OPRO CriSPO CriSPO 3-shot*
@@
ACI-Bench Claude Instant 26.1 33.5 25.5 (±1.0) 26.8 (±1.4) 35.3 (±2.3)
@@
Dataset LLM 0-shot 3-shot* OPRO CriSPO CriSPO 3-shot*
@@
ACI-Bench Claude Instant 85.5 88.1 85.1 (±0.3) 85.8 (±0.7) 88.1 (±0.5)
@@
Dataset LLM 0-shot 3-shot* OPRO CriSPO CriSPO 3-shot*
@@
ACI-Bench Claude Instant 66.7 66.3 62.3 (±2.0) 63.3 (±3.3) 65.6 (±1.1)
@@
Appendix N Standard Deviation for QA
@@
Dataset Claude 0-shot 64-shot OPRO CriSPO CriSPO 64-shot
@@
0-shot 5-shot OPRO CriSPO CriSPO 5-shot